Given that the electric field between the plates (inside the membrane) is about 107 N/C and the thickness of the membrane is about 7 nm, estimate the acceleration of the ion in nm/s2 if it were in the channel (and the channel were empty). The ion is an Na+.
Answer #1

Force on ion is F=RE F= 1.6 X1619x107 N = 1.6x10-12 N. a = 5 = 1.6x1024 0-04166 x 10 m/s2 a= 23* 1.6 7x102Fkg a = 41.66x10m s